OpenStudy (goformit100):

How to do it mentally ? Peter has 20 dollars in his pocket. He buys a 32 ounces vanilla ice cream for 8 dollars. How much is Peter's change?

OpenStudy (anonymous):

20 - 8 = 12

OpenStudy (goformit100):

Madam can you please explain me .

OpenStudy (anonymous):

It's subtraction, how do you explain that?? -_- Hmm, okay, well I do it so you think about it, like 10 - 8 = 2, and then you still have 10, 10 + 2 = 12. Thats the best way I can explain this -.-

OpenStudy (goformit100):

Thankyou madam